Assume the following cost information for Fernandez​ Company: Selling price $180 per unit Variable costs $60 per unit Total fixed costs $85,000 Tax rate 40​% What minimum volume of sales dollars is required to earn an afterminus−tax net income of $60,000​? ​(Do not round interim calculations and round the final answer to the nearest​ dollar.) A. $277,500 B. $150,000 C. $212,500 D. $127,500
Answer #1

Answer is A: $ 277500

Let the selling units =x

Than,

Sales-variable cost- fixed cost= before tax profit

Before tax profit= $ 60000/40%=$100000

So,

(180-60)x-85000=100000

120x-85000=100000

120x=185000

x=1541.66666667 units

Sales in dollar =1541.6666667 *180

= $ 277500